Kyra's basis in investment use land was $9000. She received a gross condemnation award of $13000. Her net condemnation award was $12600. She purchased replacement property at a cost of $13000. What is Kyra's basis in the replacement property?

s
Answer #1

Unrealized gain = Net condemnation award - Basis in investment

= 12,600 - 9,000 = 3,600

Basis in the replacement property

= Cost of replacement property purchased - Unrealized gain

= 13,000 - 3,600

= 9400